# Optical-field-induced dips and splits in nonlinear spectra of selective reflection from high-density atomic vapor

### v1

@@ Abstract
We discuss nonlinear spectra of selective reflection from high-density rubidium atomic vapor, where the self-broadening of the resonant transition $5S_{1/2}-5P_{3/2}$ dominates over the Doppler width. In the experiments, the hole-burning technique with probe and pump lasers is used. The reflection of weak probe beam is investigated at four atomic densities in the range $(1.2\text{--}3.6)\cdot 10^{17}$ cm-3 and various pump beam intensities. To enhance the spectral resolution the frequency derivative $\text{d}R/\text{d}\nu$ of the reflection coefficient $R$ recorded and analyzed. The grow of the atomic number density changes the character of self-broadening from inhomogeneous to homogeneous. At the maximal density, the strong pump field split observed spectra on two homogeneously broadened symmetrical resonances. The appearance of the optical field induced resonances can be explained in the frame of ”dressed atomic states” approach. In the range of the lower densities the spectral profiles are in-homogeneously broadened. Spectral profiles of the frequency derivative are separated by the optically saturated dips. The width of such dips is a superposition of the homogeneous part of self-broadening and intensity dependent field broadening. Careful study of the transition from inhomogeneous to homogeneous broadening may be initiate a further development of theory of interatomic interactions in high density atomic gas media.
@@ 1 Introduction
In high-density atomic gas, where the dipole self-broadening of resonance transitions is much larger than the Doppler width, the thermal motion of atoms can be neglected Lewis1980 , BoydPRL1991 . The self-broadening $\Gamma_{0}$ in such undisturbed gas of atomic number density $N$ can be evaluated as $\Gamma_{0}=KN$, where $K\sim 10^{-16}$ GHz cm3. The spectral properties of the high-density atomic gases can be recast in the framework of disordered excitons MukamelPRA1994 . Static dipole–dipole interactions between atoms can induce a combination of homogeneous and inhomogeneous broadening of atomic transitions. Possible many-body effects should also be considered. Homogeneous or inhomogeneous character of the line shapes may be tested by applying nonlinear laser spectroscopy methods such as photon echoes or hole burning MukamelJCP1991 . The theoretical publications MukamelPRA1994 , MukamelJCP1991 motivated to conduct an experimental research of the nonlinear selective reflection from the high-density atomic vapor VASPRL1996 . In the experiment with rubidium vapor the probe and pump tunable lasers were used. The probe laser frequency was scanned over the $5S_{1/2}-5P_{3/2}$ transition. The pump laser frequency was fixed in the far wing of the atomic transition, where the absorption length was much longer than the wavelength. The off-resonant optical excitation of rubidium atoms is incoherent due to radiation trapping and interatomic excitation transfer HolsteinPRA1947 , Biberman1947 , VASJQSRT2020 . As a result of a comparison of experimental and calculated reflection spectra a linear relation between self-broadening $\Gamma$ and the ground state population $N_{\text{g}}$ VASPRL1996 is derived $\Gamma=KN_{\text{g}}$.
@@
The observed excitation dependence is explained by static mechanism of the dipole-dipole interactions in the high-density atomic vapor MukamelPRA1994 . Dipole-dipole interactions appeared only between the ground and excited atoms. Subsequent measurements VASPRA1999 , VASJPhysB2009 confirmed the derived linear relation for self-broadening in incoherently excited potassium and rubidium atomic vapors.
@@
The optical resonance saturation of high-density rubidium vapors is investigated in selective reflection experiments using one VASJRLR2021 , BobrovJETPLett2021 , VASJQSRT2022 and two tunable lasers VASNSRJQSRT2024 . Intense laser emission induces a reduction in the width and magnitude of the selective reflection spectral profiles. Field-broadening effects are also discussed.
@@
Interesting nonlinear phenomena may be observed in coherently excited resonant gas, when the Rabi oscillations dominate over relaxation processes Demtroderbook2008 , Boydbook2007 . Observation of the coherent Autler–Townes effect is reported in the seminal publication AutlerTownes1955 . The Autler–Towns components were observed in selective reflection from the high-density potassium vapor VASATPRA2008 . The probe laser frequency was scanned over the $4S_{1/2}-4P_{1/2}$ transition. The pump laser frequency was fixed at the $4S_{1/2}-4P_{3/2}$ transition. The appearance of the Autler–Townes doublet is interpreted within the framework of ”dressed atomic states” approach CohenTannoudjiATbook1996 , CohenTannoudjibook1992 . The measured spectral splitting of the ground state $4S_{1/2}$ is defined by the Rabi frequency.
@@
It has been shown theoretically KaiserPRA2019 that the ”dressed dense atomic gases” may be used to explore the many-body effects. The main goal of our current research is to identify experimental conditions for the formation of the ”dressed states” in the high-density atomic vapors. The optical preparation of new atomic levels can be described by a simple semiclassical model Demtroderbook2008 . In this framework, the optical coherent excitation of an ensemble of two-level atoms with an unperturbed transition frequency $\omega_{ab}$, a transition dipole moment $D_{ab}$ and a homogeneous width $\gamma$ (FWHM) is considered. The excitation rate of atoms, when the frequency of the applied optical field with amplitude $E_{0}$ is tuned at the frequency $\omega_{ab}$, is defined by the ratio of the optically saturated width $\gamma_{\text{sat}}/2$ and the resonance Rabi frequency $\Omega_{\text{R}}$ Demtroderbook2008 :
@@
When $\Omega_{\text{R}}$ exceeds $\gamma_{\text{sat}}/2$, the Rabi oscillations lead to formation of sidebands at frequencies $\omega_{ab}\pm\Omega_{\text{R}}$. In the absorption spectrum, the symmetrical resonances with a saturated width of $\gamma_{\text{sat}}$ will appear at these frequencies. The frequency interval between the absorption resonances $\Delta\omega_{\text{abs}}$ equals to $2\Omega_{\text{R}}$. According the ”dressed atomic states” approach the strong optical field induces new energy levels by dynamic Stark shifts CohenTannoudjibook1992 , Boydbook2007 .
@@
Analysis of selective-reflection spectra is more complicated than analysis of absorption spectra. The spectral profile of unsaturated selective reflection can be roughly approximated by a dispersive curve with a width of $\Gamma_{0}$ BoydPRL1991 , VASPRA1999 . The spectral resolution in this case is limited by slowly decaying wings: $R\sim(\omega-\omega_{ab})$. The derivative $\text{d}R/\text{d}\omega$ gives a bell-shaped spectral profile (absorption-like resonance) with rapidly decreasing wings $R\sim(\omega-\omega_{ab})^{-2}$. Using this technique, the spectral resolution in the selective reflection experiments can be noticeably improved VASATPRA2008 . The resulting spectral curves can be analyzed by using the description of the theoretical absorption spectra Demtroderbook2008 .
In the presented study the derivative of the selective reflection coefficient $\text{d}R/\text{d}\nu$ (frequency $\nu=\omega/2\pi$) are recorded and analyzed at several atomic number densities and varying saturating laser beam intensities.
Figure 2: Spectral dependence of normalized reflection coefficient $(R-R_{0})/R0$ for the probe beam, measured at four number densities: (a) $N_{1}=1.2\times 10^{17}$ cm-3 ($\Gamma_{0}/2\pi=13.2$ GHz), (b) $N_{2}=1.7\times 10^{17}$ cm-3 ($\Gamma_{0}/2\pi=18.7$ GHz), (c) $N_{3}=2.5\times 10^{17}$ cm-3 ($\Gamma_{0}/2\pi=27.5$ GHz), and (d) $N_{4}=3.6\times 10^{17}$ cm-3 ($\Gamma_{0}/2\pi=39.6$ GHz).
Figure 3: Frequency derivatives of the reflection coefficient $\text{d}R/\text{d}\nu$ at four different number densities $N_{1}$, $N_{2}$, $N_{3}$, $N_{4}$ and three selected values of the pump beam intensity $I_{\text{pump}}=0$ kW/cm2 (a), $I_{\text{pump}}=4.8$ kW/cm2 (b), $I_{\text{pump}}=8.8$ kW/cm2 (c). The bottom curves show saturated-absorption spectra of the probe beam for the reference rubidium cell. The frequency scale is the same as in Fig. 2.
@@ 2 Experimental setup
Nonlinear selective reflection from the window–vapor interface in the high-temperature vapor cell with natural abundance of 85Rb and 87Rb isotopes was studied at the transition $5S_{1/2}-5P_{3/2}$ (780 nm). The transition consists hyperfine (hfs) components of $^{85}Rb$ and 87Rb isotopes. A brief description of the used vapor cell is given in VASselffoc2019 . The experiments are conducted at four atomic number densities $N_{i}$ in the range $(1.2\text{--}3.6)\cdot 10^{17}$ cm-3 which defined by the temperatures of the cell from $367$ to $427^{\circ}$C. Self-broadening $\Gamma_{0}=KN$ of the working transition is estimated by using the factor $K/2\pi=(1.1\pm 0.17)\cdot 10^{-16}$ GHz cm3 from AdamsJPhysB2011 .
@@
The setup is practically the same as it used in VASNSRJQSRT2024 . Spectroscopic measurements are performed using two tunable diode lasers: a low-power probe laser and a high-power pump laser. The probe lasers frequency is scanned over all hfs-components of $5S_{1/2}-5P_{3/2}$ transition. To calibrate the probe lasers frequency $\nu_{\text{pr}}$, we record the saturated absorption spectrum of the reference rubidium vapor cell with natural abundance of 85Rb and 87Rb isotopes at a room temperature. In our research the hfs-component $5S_{1/2}(F=3)-5P_{3/2}(F^{\prime}=4)$ in 85Rb atoms is chosen as the frequency reference $\nu_{0}$. The pump lasers frequency $\nu_{\text{pump}}$ is fixed at the reference $\nu_{0}$. Schematic of our setup is presented in Fig. 1.
@@
The pump and probe optical beams with orthogonal linear polarization are combined on polarizing beam splitter (PBS) and then focused by lens L onto the inner surface of the cells input window. Application of the pump and probe beams with orthogonal polarizations allows to eliminate possible four-wave mixing at the window/rubidium vapor interface VASFWMPRA1997 , CASFWMQuEl2023 . The pump beam intensity can be varied from zero to 8.8 kW cm-2. The probe beam intensity $I+\text{pr}$ is kept below 4 W cm-2 to avoid a distortion of the recorded curves which could be induced by the probe beam. The probe beam is incident at the 64 mrad. The angle of the incidence for the pump beam is kept at 4 mrad. Probe beam reflected from the interface widow/vapor is directed to a PD photodetector. Scattered radiation from the saturating beam is additionally suppressed by a GT polarizer. The reflection signal from the PD photodetector is recorded using a digital oscilloscope and a computer. The measured selective reflection coefficient $R$ is normalized by the non-resonant reflection coefficient $R_{0}\approx 8.5$% from the window/vacuum interface as $\delta R=(R-R_{0})/R_{0}$. To ensure the higher spectral resolution, we record the derivative of the selective reflection coefficient $\text{d}R/\text{d}\nu$ by computer processing of the experimental data.
@@ 3 Experimental results and discussions
The spectral dependencies of the normalized reflection coefficient $\delta R$, recorded at different atomic number densities and pump intensities, are shown in Fig. 2. The frequency scale is expressed as detuning $\Delta_{\text{pr}}/2\pi=(\nu_{\text{pr}}-\nu_{0})$ of the probe laser from the hfs component $5S_{1/2}(F=3)-5P_{3/2}(F^{\prime}=4)$. The pump laser frequency $\nu_{\text{pump}}$ was fixed at the reference frequency $\nu_{0}$. The resonance optical saturation modifies the spectral profiles of $\delta R$, including splitting of the spectra. The width of the unsaturated transition $5S_{1/2}-5P_{3/2}$, which is a combination of self-broadening and hfs-splitting, can be estimated as the frequency interval between minimum and maximum of the $\delta R$ profiles in absence of the pump beam BoydPRL1991 , VASPRA1999 . More reliable results can be reached by measuring the frequency intervals $\Delta\nu_{\text{SR}}$ between ”zeros” of the $\text{d}R/\text{d}\nu$ function VASPRA1999 . The recorded derivatives of the unsaturated reflection coefficient $\text{d}R/\text{d}\nu$ are presented in Fig. 3(a).
According to MukamelPRA1994 in the high-density atomic gas the dipole self-broadening can be a combination of homogeneous and inhomogeneous widths. The hole burning technique can help to estimate contributions of the homogeneous and inhomogeneous parts to the linewidth. In the Figs. 3(b) and 3(c) the nonlinear spectra of frequency derivative $\text{d}R/\text{d}\nu$ are presented. The dips at the pump laser frequency separate the spectra of $\text{d}R/\text{d}\nu$ on two resonances. At densities $N_{1}$, $N_{2}$ and $N_{3}$ the resonances are asymmetrical. The spectral difference between resonances can be attributed to the influence of the hfs-components. Such properties of the spectra indicate that the resonances at densities $N_{1}$, $N_{2}$ and $N_{3}$ are inhomogeneously broadened. Dips are results of the optical saturation of the resonance transitions Demtroderbook2008 . The width of each dip is a superposition of the homogenous part of self-broadening and field broadening Demtroderbook2008 .
@@
Curves in Figs. 3(b), 3(c), which recorded at the maximal density $N_{4}$, demonstrate splitting of spectra of $\text{d}R/\text{d}\nu$ on two symmetrical resonances. It means that the symmetrical resonances are homogeneously broadened. The spectral properties of the symmetrical resonances can be explained by Rabi flopping oscillations in the simple semiclassical model Demtroderbook2008 or by dynamic Stark shifts in the ”dressed states” description CohenTannoudjibook1992 , Boydbook2007 . In both representations the frequency interval between symmetrical resonances is defined by the Rabi frequency. We shall note that a definition of the resonance Rabi frequency $\Omega_{R}$ in Demtroderbook2008 and a definition of the Rabi frequency $\Omega$ in CohenTannoudjibook1992 , Boydbook2007 differ by factor two, as result, $\Omega=2\Omega_{R}$. For the working transition $5S_{1/2}-5P_{3/2}$ in rubidium atoms, the frequency interval $\Delta\nu=2\Omega_{R}$ can be expressed as a function of the pump intensity $I_{\text{pump}}$ VASJRLR2021 :
$\Delta\nu~\text{[GHz]}=2.53\cdot(I_{\text{pump}}\text{[kW/cm}^{-2}\text{]})^{1/2}.$ (2)
@@
The measurements of the splitting interval $\Delta\nu_{\text{SP}}$ between the symmetrical resonances confirm our suggestion, that new levels can be explained by ”atomic dressed states” description, are showed up in our experiment. The experimental data and fitting lines are presented in Fig. 4.
@@
In the Table 1 the fitted parameters $a$ and $b$ are presented for the different atomic densities $N_{i}$.

### v2

@@ Abstract
We discuss nonlinear spectra of selective reflection from high-density rubidium atomic vapor, where the self-broadening of the resonant transition $5S_{1/2}-5P_{3/2}$ dominates over the Doppler width. In the experiments, the hole-burning technique with probe and pump lasers is used. The reflection of weak probe beam is investigated at four atomic densities in the range $(1.2\text{--}3.6)\times 10^{17}$ cm-3 and various pump beam intensities. To enhance the spectral resolution, the frequency derivative $\text{d}R/\text{d}\nu$ of the reflection coefficient $R$ is analyzed. Increasing the atomic number density changes the character of self-broadening from inhomogeneous to homogeneous. At the highest density, the strong pump field splits the observed spectra into two homogeneously broadened symmetric resonances. The appearance of the optical-field-induced resonances can be explained within the framework of ”dressed atomic states” approach. At lower densities the spectral profiles are inhomogeneously broadened. Spectral profiles of the frequency derivative are separated by optically saturated dips. The width of such dips is a combination of the homogeneous component of self-broadening and intensity-dependent field broadening. Careful study of the transition from inhomogeneous to homogeneous broadening may initiate further development of the theory of interatomic interactions in high density atomic gas media.
@@ 1 Introduction
In high-density atomic gas, where dipole-dipole self-broadening of the resonance transitions is much larger than the Doppler width, the thermal motion of atoms can be neglectedLewis1980, BoydPRL1991. The self-broadening $\Gamma_{0}$ in such gas of atomic density $N$ can be estimated by using a relation $\Gamma_{0}=KN$, where factor $K\sim 10^{-16}$ GHz cm3. The theory of self-broadening was extended in the framework of disordered excitons in Ref. MukamelPRA1994. The developed approach was applied to the study of dense two-level atomic gases. The main part of the Hamiltonian used described dipole-dipole interactions between the ground $S$-state and excited $P$-state atoms. Possible contributions of interactions between atoms in the same quantum states were neglected. It was noted that spectral profiles of resonance transitions were usually classified as either homogeneous or inhomogeneous. A spectral line shape was considered to be homogeneously broadened when the dominant broadening mechanism was due to the binary collisions of the moving atoms (collisional broadening), and inhomogeneous if the spectral width was mainly caused by a distribution of transition frequencies (static broadening). The inhomogeneous self-broadening induced by many-particle interactions was explained by the developed model of frequency shifted excitons with long-range interactions. The concept of the Lorentz local field in a disordered medium was also used. In the realm of linear optics, it is impossible to distinguish self-broadened line shapes in a dense atomic gas. In the discussed approach both broadening mechanisms gave Lorentzian-like spectral profiles and, moreover, the calculated static and collision widths are of comparable magnitudes, with a fixed ratio of 3:8, regardless of the atomic density.
Notably, there must be a threshold density above which the binary collisions become interrupted by third particles. Such an event would occur when the average interparticle distance becomes of less or the same order of magnitude as the Weisskopf radius. A density $N_{\mathrm{W}}$, which defined by Weisskopf radius for potassium atomic vapor, was estimated to be $10^{18}$ cm-3. At densities $N\geq N_{\mathrm{W}}$, the dipole-dipole interactions are altered. The discussed theoretical treatment MukamelPRA1994 was performed for much less densities, $N\ll N_{\mathrm{W}}$. Under these conditions the short-range deviations from the dipole-dipole interactions (influence of higher multipoles) were neglected. Predicted homogeneous or inhomogeneous character of the spectral profile of atomic transition could be tested by nonlinear laser spectroscopy methods such as photon echoes or hole burning MukamelJCP1991. The publications MukamelPRA1994, MukamelJCP1991 motivated authors of the Ref. VASPRL1996 to perform an experimental study of a nonlinear selective reflection from the dense rubidium atomic vapor.
@@
In the experiment on rubidium vapor VASPRL1996 the probe and pump tunable lasers were used. The probe laser frequency was scanned over the $5S_{1/2}-5P_{3/2}$ transition. The pump laser frequency was fixed in the far wing of the atomic transition, where the absorption length was much longer than the wavelength. The off-resonant optical excitation of rubidium atoms was incoherent due to radiation trapping and interatomic excitation transfer HolsteinPRA1947, Biberman1947, VASJQSRT2020. As a result of comparison of experimental and calculated reflection spectra a linear relation between self-broadening $\Gamma$ and the ground state population $N_{\text{g}}$ VASPRL1996 is derived $\Gamma=KN_{\text{g}}$.
@@
The observed excitation dependence is explained by the static mechanism of dipole-dipole interactions in the high-density atomic vapor MukamelPRA1994. Dipole-dipole interactions appeared only between the ground and excited atoms. Subsequent measurements VASPRA1999, VASJPhysB2009 confirmed the derived linear relation for self-broadening ($\Gamma=KN_{\text{g}}$) in incoherently excited potassium and rubidium atomic vapors.
@@
The optical resonance saturation of high-density rubidium vapors is investigated in selective reflection experiments using one VASJRLR2021, BobrovJETPLett2021, VASJQSRT2022 and two tunable lasers VASNSRJQSRT2024. Intense laser emission reduces both the width and magnitude of the selective reflection spectral profiles. Field-broadening effects are also discussed.
@@
Interesting nonlinear phenomena may be observed in coherently excited resonant gas, when the Rabi oscillations dominate over relaxation processes Demtroderbook2008, Boydbook2007. Observation of the coherent Autler–Townes effect is reported in the seminal publication AutlerTownes1955. The Autler–Townes components were observed in selective reflection from the high-density potassium vapor VASATPRA2008. The probe laser frequency was scanned over the $4S_{1/2}-4P_{1/2}$ transition. The pump laser frequency was fixed at the $4S_{1/2}-4P_{3/2}$ transition. The appearance of the Autler–Townes doublet was interpreted within the framework of ”dressed atomic states” approach CohenTannoudjiATbook1996, CohenTannoudjibook1992. The measured spectral splitting of the ground state $4S_{1/2}$ was defined by the Rabi frequency.
@@
It has been shown theoretically KaiserPRA2019 that the ”dressed dense atomic gases” may be used to explore the many-body effects. The main goal of our current research is to identify experimental conditions for the formation of the ”dressed states” in the high-density atomic vapors. The optical preparation of new atomic levels can be described by a simple semiclassical model Demtroderbook2008. In this framework, the optical coherent excitation of an ensemble of two-level atoms with an unperturbed transition frequency $\omega_{ab}$, a transition dipole moment $D_{ab}$ and a homogeneous width $\gamma$ (FWHM) was considered. When the applied optical field of amplitude $E_{0}$ was tuned at the frequency $\omega_{ab}$, the atomic excitation rate was given by the ratio of the optically saturated width $\gamma_{\text{sat}}/2$ and the resonance Rabi frequency $\Omega_{\text{R}}$ Demtroderbook2008:
@@
For $\Omega_{\text{R}}>\gamma_{\text{sat}}/2$, Rabi oscillations lead to formation of sidebands at frequencies $\omega_{ab}\pm\Omega_{\text{R}}$. In the absorption spectrum, the symmetric resonances with a saturated width of $\gamma_{\text{sat}}$ will appear at these frequencies. The frequency interval $\Delta\omega_{\text{abs}}$ between the absorption resonances is equal to $2\Omega_{\text{R}}$. According to the ”dressed atomic states” approach the strong optical field induces new energy levels by dynamic Stark shifts CohenTannoudjibook1992, Boydbook2007.
@@
Analysis of selective-reflection spectra is more complicated than analysis of absorption spectra. The spectral profile of unsaturated selective reflection can roughly be approximated by a dispersive curve with a width of $\Gamma_{0}$ BoydPRL1991, VASPRA1999. The spectral resolution in this case is limited by slowly decaying wings: $R\propto(\omega-\omega_{ab})^{-1}$. The derivative $\text{d}R/\text{d}\omega$ gives a bell-shaped spectral profile (absorption-like resonance) with rapidly decreasing wings $\text{d}R/\text{d}\omega\propto(\omega-\omega_{ab})^{-2}$. Using this technique, the spectral resolution in the selective reflection experiments can be substantially improved VASATPRA2008. The resulting spectral curves can be analyzed by comparison with the theoretical absorption spectra Demtroderbook2008.
In the presented study the derivative $\text{d}R/\text{d}\nu$ ($\nu=\omega/2\pi$) of the selective reflection coefficient $R$ was analyzed at several atomic densities and pump laser beam intensities. We take the derivative numerically from the spectral dependence of $R$.
Figure 2: Spectral dependence of normalized reflection coefficient $(R-R_{0})/R_{0}$ for the probe beam, measured at four number densities: (a) $N_{1}=1.2\times 10^{17}$ cm-3 ($\Gamma_{0}/2\pi=13.2$ GHz), (b) $N_{2}=1.7\times 10^{17}$ cm-3 ($\Gamma_{0}/2\pi=18.7$ GHz), (c) $N_{3}=2.5\times 10^{17}$ cm-3 ($\Gamma_{0}/2\pi=27.5$ GHz), and (d) $N_{4}=3.6\times 10^{17}$ cm-3 ($\Gamma_{0}/2\pi=39.6$ GHz). Zero frequency corresponds to the transition $5S_{1/2}(F=3)-5P_{3/2}(F^{\prime}=4)$ in 85Rb atoms.
Figure 3: Frequency derivatives of the reflection coefficient $\text{d}R/\text{d}\nu$ at four different number densities $N_{1}$, $N_{2}$, $N_{3}$, $N_{4}$ and three selected values of the pump beam intensity (a) $I_{\text{pump}}=0$ kW/cm2, (b) $I_{\text{pump}}=4.8$ kW/cm2, (c) $I_{\text{pump}}=8.8$ kW/cm2. The bottom curves show saturated-absorption spectra of the probe beam for the reference rubidium cell. The frequency scale in panel (a) is the same as in Fig. 2. Zero frequency corresponds to the transition $5S_{1/2}(F=3)-5P_{3/2}(F^{\prime}=4)$ in 85Rb atoms. Horizontal dashed lines show the ”zero” levels. Vertical dotted lines indicate the location of the pump laser frequencies. Marks (x) indicate crossing of the derivative curves with the ”zero” levels (dashed lines).
@@ 2 Experimental setup
Nonlinear selective reflection from the window–vapor interface in the high-temperature vapor cell with natural abundance of 85Rb and 87Rb isotopes was studied at the transition $5S_{1/2}-5P_{3/2}$ (780 nm). This transition consists of two hyperfine (hfs) components from 85Rb and 87Rb isotopes. A brief description of the used vapor cell is given in VASselffoc2019. The experiments are conducted at four atomic number densities $N_{i}$ in the range $(1.2\text{--}3.6)\times 10^{17}$ cm-3 which defined by the temperatures of the cell from $367$ to $427^{\circ}$C. Self-broadening $\Gamma_{0}=KN$ of the working transition is estimated by using the factor $K/2\pi=(1.1\pm 0.17)\times 10^{-16}$ GHz cm3 from AdamsJPhysB2011.
@@
The setup is practically the same as was used in Ref. VASNSRJQSRT2024. Spectroscopic measurements are performed using two tunable diode lasers: a low-power probe laser and a high-power pump laser. The probe laser frequency is scanned over all hfs components of $5S_{1/2}-5P_{3/2}$ transition. To calibrate the probe laser frequency $\nu_{\text{pr}}$, we record the saturated absorption spectrum of the reference rubidium vapor cell with natural abundance of 85Rb and 87Rb isotopes at room temperature. In our measurements, the hyperfine (hfs) component $5S_{1/2}(F=3)-5P_{3/2}(F^{\prime}=4)$ in 85Rb atoms is chosen as the frequency reference $\nu_{0}$. The pump laser frequency $\nu_{\text{pump}}$ is fixed at the reference $\nu_{0}$. A schematic of the setup is shown in Fig. 1.
@@
The probe and pump beams with orthogonal linear polarization are combined by a polarizing beam splitter and then focused onto the window–vapor interface in the cell. The orthogonal polarizations are used to prevent possible formation of narrow resonances due to the coherent scattering of probe and pump optical fields by induced oscillations of atomic states population at the beat frequency VASFWMPRA1997. The contrast coherent resonances in a spectral dependence of the reflected probe beam were studied by using two beams with the same polarization CASFWMQuEl2023. The selective reflection from the transparent dielectric–disordered gas interface is practically independent on a polarization of the incident optical beam near the normal angle of incidence. The pump beam intensity can be varied from zero to 8.8 kW cm-2. The probe beam intensity $I_{\text{pr}}$ is kept below 4 W cm-2 to avoid a distortion of the recorded curves which could be induced by the probe beam. The probe beam is incident at 64 mrad. The angle of incidence for the pump beam is kept at 4 mrad. The probe beam reflected from the interface window/vapor is directed to a photodiode (PD). Scattered radiation from the pump beam is additionally suppressed by a Glan-Thompson polarizer (GT). The presence of the GT polarizer enhances contrast and does not change the measured selective reflection coefficient. The selective reflection coefficient was calibrated using the cell at room temperature and at large detuning of the pump beam frequency. The reflection signal from the PD photodiode is recorded with a digital oscilloscope and a computer. The measured selective reflection coefficient $R$ is normalized by the nonresonant reflection coefficient $R_{0}\approx 8.5$% from the window/vacuum interface as $\delta R=(R-R_{0})/R_{0}$. To ensure higher spectral resolution, we recorded the selective-reflection spectrum $R$ and averaged it using a digital oscilloscope to improve the signal-to-noise ratio. The frequency derivative $\text{d}R/\text{d}\nu$ was then calculated numerically during data processing. The dispersive selective-reflection profile has slowly decaying wings, whereas its frequency derivative is bell-shaped with rapidly decreasing wings, thereby allowing closely spaced spectral features to be resolved more clearly.
@@ 3 Experimental results and discussions
The spectral dependencies of the normalized reflection coefficient $\delta R$, recorded at different atomic number densities and pump intensities, are shown in Fig. 2. The frequency scale is expressed as detuning $\Delta_{\text{pr}}/2\pi=(\nu_{\text{pr}}-\nu_{0})$ of the probe laser from the hfs component $5S_{1/2}(F=3)-5P_{3/2}(F^{\prime}=4)$. The pump laser frequency $\nu_{\text{pump}}$ was fixed at the reference frequency $\nu_{0}$. The resonance optical saturation modifies the $\delta R$ spectral profiles, leading to including splitting of the spectra. The width of the unsaturated transition $5S_{1/2}-5P_{3/2}$, which is a combination of self-broadening and hfs-splitting, can be estimated as the frequency interval between minimum and maximum of the $\delta R$ profiles in the absence of the pump beam BoydPRL1991, VASPRA1999. More reliable results can be reached by measuring the frequency intervals $\Delta\nu_{\text{SR}}$ between ”zeros” of the $\text{d}R/\text{d}\nu$ function VASPRA1999. Calculated derivatives of the unsaturated reflection coefficient $\text{d}R/\text{d}\nu$ are presented in Fig. 3(a). The marks (x) indicate crossing points of the unsaturated spectral dependence of derivative with the ”zero” level (dashed line) in Fig. 3(a). According to MukamelPRA1994 in the high-density atomic gas the dipole self-broadening can be represented as a combination of homogeneous and inhomogeneous widths. The hole-burning technique can help to estimate contributions of the homogeneous and inhomogeneous parts to the linewidth. In the Figs. 3(b) and 3(c) the nonlinear spectra of frequency derivative $\text{d}R/\text{d}\nu$ are presented. The dips at the pump laser frequency split the spectra of $\text{d}R/\text{d}\nu$ into two resonances. At densities $N_{1}$, $N_{2}$ and $N_{3}$ the resonances are asymmetric. The spectral difference between resonances can be attributed to the influence of the hfs components. Such properties of the spectra indicate that the resonances at densities $N_{1}$, $N_{2}$ and $N_{3}$ are inhomogeneously broadened. Dips are the result of the optical saturation of the resonance transitions Demtroderbook2008. The width of each dip is a superposition of the homogeneous part of self-broadening and the field broadening Demtroderbook2008.
@@
Curves in Figs. 3(b), 3(c), which recorded at the maximal density $N_{4}$, demonstrate splitting of spectra of $\text{d}R/\text{d}\nu$ into two symmetric resonances. In our analysis, we consider the spectral width to be a more informative parameter than the amplitude. The double resonances with the same spectral widths can be considered as symmetric. Which implies that the symmetric resonances are homogeneously broadened. The spectral properties of the symmetric resonances can be explained either by the Rabi flopping oscillations in the simple semiclassical model Demtroderbook2008 or by the dynamic Stark shifts in the ”dressed states” approach CohenTannoudjibook1992, Boydbook2007. In both representations the frequency interval between symmetric resonances is determined by the Rabi frequency. Note that the resonance Rabi frequency $\Omega_{\text{R}}$ defined in Ref. Demtroderbook2008 and the Rabi frequency $\Omega$ used in Refs. CohenTannoudjibook1992, Boydbook2007 differ by a factor of two, such that $\Omega=2\Omega_{\text{R}}$. For the working transition $5S_{1/2}-5P_{3/2}$ in rubidium atoms, the frequency interval $\Delta\nu=2\Omega_{\text{R}}$ is related to the pump intensity $I_{\text{pump}}$ VASJRLR2021 as follows:
$\Delta\nu~\text{[GHz]}=2.53\cdot(I_{\text{pump}}\text{[kW/cm}^{2}\text{]})^{1/2}.$ (2)
@@
The measurements of the splitting interval $\Delta\nu_{\text{SP}}$ between the symmetric resonances confirm our suggestion, that new levels can be described as ”atomic dressed states”. The experimental data and fitting lines are presented in Fig. 4. As one can see on Fig. 3(a) and Fig. 3(b) for $N_{1}$ there is a small shift of the narrow symmetric resonances relative to the pump frequency. The frequency splitting is defined by the generalized Rabi frequency which depends on the resonance Rabi frequency and the frequency shift Boydbook2007. In our case the frequency shift induced contribution to the measured splitting is small relatively to experimental errors and can be neglected.
@@
In the Table 1 the fitted parameters $a_{i}$ and $b_{i}$ are presented for the different atomic densities $N_{i}$.